OpenStudy (anonymous):

What is the absolute value of -2+3i?

OpenStudy (anonymous):

The "absolute-value" or sometimes called the "modulus" of a complex number a+ bi is defined as the sqrt(a^ + b^2). So the complex number, 5-12i, will have a "modulus" or "absolute-value" of sqrt ((5^2) + (-12)^2) = sqrt (25+144) = sqrt (169) = 13.

OpenStudy (anonymous):

Thanks!!!!

OpenStudy (anonymous):

welcome.